OpenStudy (anonymous):

Identify the 15th term of the geometric sequence 52, 39, 29.25

OpenStudy (zepp):

What's the common ratio? (r)

OpenStudy (zepp):

It would be 0.75 (39/52 = 0.75 29.25/39 = 0/75) And the formula to find any term in a geometric sequence is \[\huge x_{n} = ar^{n-1}\]

OpenStudy (zepp):

The first term is 52 (a) \[\huge x_{15} = 52 * 0.75^{15-1} \approx 0.9265\]